# Zhang Q Y, Xu L, Huang J H, et al

### v1

@@ 2.3 Reconfigurable Satellite Formation Flying
Moreover, the virtual structure formation method takes a different approach by treating the formation as a single “rigid body” [36]. Each satellite maintains a fixed position within a virtual structure, such as a wheel, line, or grid. This method is effective for achieving high precision in maintaining formation shape and alignment. For example, wheel formations involve satellites distributed around a circular or ring-like path, providing 360-degree coverage ideal for tasks like surveillance and communication relays. Other configurations include linear formations, which arrange satellites in a straight line, suitable for continuous EO and synchronized orbit tracking, and rectangular or grid formations, where satellites form a structured grid pattern for broad-area coverage. Triangular formations and cubic or spherical formations are also common in virtual structure control, offering three-dimensional coverage for complex space missions that require multi-directional observations or high spatial resolution. Additionally, honeycomb formations follow a hexagonal grid structure, maximizing spatial efficiency and are often used in communication networks and environmental monitoring due to their dense, uniform coverage. Arrow or V-formations are another variant, typically used in applications that require directional alignments, such as clustered navigation or coordinated maneuvering, as they reduce aerodynamic drag (for atmospheric applications) and optimize the communication link among satellites. Beyond these conventional configurations, recent research has also introduced optimization and predictive control techniques into satellite formation control. For instance, optimal control methods aim to minimize fuel consumption or time delay, essential for deep-space missions with limited resources. Model predictive control (MPC) allows satellites to adjust their trajectories based on predictions of future system states, providing adaptive and efficient control for complex formations with high maneuverability demands.
@@
Several research efforts have focused on developing the necessary technologies to achieve RSFF. One of the most prominent approaches is the use of graph theory to model inter-satellite communication and control [40]. By representing satellites as nodes in a graph and their communication links as edges, researchers have developed algorithms to optimize formation control under various constraints, including limited communication bandwidth, time-varying network topologies, and external disturbances. Another key area of research involves the development of formation control strategies based on artificial potential functions. These methods model satellites as particles moving within a potential field generated by their relative positions and target locations. By adjusting the forces acting on each satellite, the formation can be reconfigured dynamically to achieve desired geometries while avoiding collisions or other hazards. This approach has proven effective in both simulation and experimental settings, particularly for missions involving close-proximity operations. Moreover, the control methods for swarm configurations share similarities with satellite formation control methods. However, for CCS system already in the desired configuration, frequent control task execution leads to resource wastage. Therefore, an event-triggered mechanism can be designed to apply control only when configuration errors exceed the predefined threshold. Event-triggered control can effectively reduce both communication frequency among members and energy consumption for orbit maintenance. Currently, event-triggered control has been increasingly utilized for maintaining swarm system configurations, specifically in research areas such as first-/second-order integral multi-agent systems [41], directed/undirected topologies [42], centralized/distributed triggering mechanisms [43], fixed/switched topologies [41], linear/nonlinear systems [44], with/without time delays [45], and input saturation [46].
@@
Further, developing decentralized control architectures that distribute decision-making across the DSIN is essential to improve resilience and flexibility. The DSIN should incorporate adaptive algorithms capable of responding to environmental changes and communication delays, enabling autonomous reconfiguration without relying on a central command unit. Moreover, hybrid centralized and decentralized control architectures use centralized control for high-precision tasks and decentralized strategies for adaptive and resilient reconfiguration, enabling the balance between precision and robustness. For inter-satellite distances exceeding 20 km, orbit maneuver requirements can be addressed based on the Lambert problem. The Lambert problem is a typical two-point boundary value problem that determines the maneuver trajectory given the spacecraft’s initial and final positions and the transfer time [47]. Considering the univariate nature of the Lambert problem, recent studies have derived transfer time equations using parameters such as flight path angle [48], eccentricity [49], and terminal velocity [50], and proposed iterative solution methods. Furthermore, maneuver theory has been improved by incorporating factors such as environmental perturbations [51], thrust direction constraints [52], and multi-revolution orbit solutions [53]. The last but not the least, researching algorithms that optimize energy consumption during formation reconfiguration is also crucial. These algorithms could focus on minimizing fuel use while considering the long-term sustainability of the mission. Techniques such as cooperative energy management and predictive energy budgeting can be explored to extend the operational lifespan of satellite constellations.
@@
By adopting a phased configuration control method at different spatial scales, the CCS systems can be restructured according to service demands, choosing system configurations and control maintenance strategies that best fit the task requirements. This enables the design of highly flexible, scalable, and open network architectures for DSIN, along with corresponding functional partitioning strategies. The future of RSFF lies in the continued integration of advanced control algorithms, high-precision sensors, and scalable communication networks. One promising avenue is the application of machine learning techniques to predict and adapt to environmental disturbances in real time. By incorporating predictive models into the control loop, satellite formations can proactively adjust their configurations to maintain optimal performance. Another important direction is adopting a phased configuration control method at different spatial scales, which allows the CCS system to be restructured according to service demands and to choose system configurations and control maintenance strategies that best fit the task requirements. This enables the design of highly flexible, scalable, and open network architectures for DSIN, along with corresponding functional partitioning strategies.
@@ 3 Enabling technologies
In order to achieve spectrum- and energy-efficient communication, as well as high timeliness and reliability, a variety of novel air interface and transmission technologies will be used in the DSIN, especially in the CCS system. In this section, we will first introduce the air interface technologies at the physical and link layers, including channel modeling and estimation, new waveforms, distributed antennas, channel coding methods, multiple access approaches, and multicasting mechanisms, followed by network and transport layer technologies, involving erasure transfer protocols, distributed routing, and congestion control.
@@ 3.1 Channel Modeling and Estimation for CCS System
As the dawn of 6G communication technologies beckons, the crucial role of DSIN in achieving uninterrupted global connectivity has become more pronounced than ever [54]. Within this evolving landscape, the creation of precise channel models and the enhancement of channel estimation methods have become pivotal to the advancement of these systems. Channel models are indispensable for capturing the intricacies of satellite-ground communication links, offering a fundamental insight into their performance across diverse operational scenarios. Concurrently, channel estimation stands as a cornerstone technology, facilitating the real-time acquisition of channel state information, which is indispensable for fine-tuning signal transmission and guaranteeing the efficiency and dependability of DSIN [55]. Despite the extensive research into terrestrial communication channel modeling and estimation, the unique challenges posed by satellite-to-ground channels in CCS system demand a closer look. In this subsection, we delve into the channel modeling and estimation within the realm of CCS system, and explore the latest advancements and present a comprehensive review of the current research landscape.
@@ 3.1.1 Channel Model
In CCS system, the channel model is essential for understanding and predicting the behavior of signals propagating between satellite-to-satellite and satellite-to-user links. We delve into the two primary categories of channel model, satellite-to-satellite and user-satellite channel model, focusing on the user-satellite channel model [56].
Satellite-to-satellite channel model, which concerns the communication links between satellites, is relatively straightforward compared to user-satellite modeling. This simplicity arises because space-to-space communication does not involve the complexities of the Earth’s atmosphere or terrestrial obstacles. As shown in Figure 7, in satellite-to-satellite channel models, the primary considerations include the relative positions and movements of the satellites, the distances between them, and the characteristics of the transmission medium, which is typically the vacuum of space. The main challenges in the satellite-to-satellite channel model involve accounting for the dynamic nature of CCS system, including the effects of gravitational forces, satellite orbit perturbations, and the need for precise alignment of communication beams. However, the relative ease of this model allows researchers to focus on optimizing the performance of ISL without the added complexities of atmospheric and terrestrial interference. The user-satellite channel model is where the real challenge lies. This model type ensures reliable communication between satellites and ground stations [57]. The user-satellite channel model is more challenging than the satellite-satellite channel model, primarily due to many factors, including the influence of the Earth’s atmosphere, interference from terrestrial obstacles, and the dynamic nature of the communication environment. User-satellite channel model methods mainly include geometric stochastic models and machine learning (ML)-based models.
@@
The geometric stochastic model is designed with rational assumptions regarding the distribution of obstacles and scatterers near the receiving user. It abstracts the effective scatterers within the environment into one or more models characterized by distinct geometric forms, systematically distributing them across the environment’s abstracted geometric structure. In [58], the studies undertook the channel model of a multi-satellite communication system, postulating that the satellites are uniformly dispersed across the surface of a sphere at a specified altitude above the Earth. By leveraging the geometric relationship between the satellite and the Earth, the research deduced the correlations among the elevation angle, azimuthal angle, and the distance between the satellite and the Earth. Although the geometric stochastic model has high universality and low computational complexity, it has low modeling accuracy in specific scenarios. To have a more accurate channel model of CCS system, much research is needed to consider more factors.
In CCS wireless communication environments, the channel conditions are more complex, the operating frequency bands are higher, and the mobility of the terminals is further enhanced. In such scenarios, ML’s self-learning and predictive capabilities are particularly suitable, as they can extract critical features from highly complex channel data. Moreover, by thoroughly training the channel model, we can enable the model to better adapt to the rapid changes in channel states in high-dynamic scenarios, thereby significantly improving the generalization ability and performance of the model. The authors in [59] integrate ML with traditional channel modeling techniques, presenting a novel approach for channel modeling in LEO satellite systems. It underscores the critical role of radio channel forecasting in improving link quality in the face of growing atmospheric impairments. In recent years, machine learning-based satellite-Earth channel modeling has seen rapid advancement, with many models being proposed, showcasing the technology’s robust learning capacity and proficiency in handling vast datasets. The LSTM-based modeling has garnered significant attention, with superior prediction accuracy demonstrated in several studies. However, the predictive accuracy of current ML-based models is contingent upon the richness of the training dataset and demands substantial computational resources, which limits the applicability of many models to specific scenarios or frequency bands, as noted in [60].
@@ 3.1.2 Channel Estimation
Channel estimation is an essential aspect of CCS systems, and it is a critical issue that directly affects the performance and reliability of the communication link. This entails estimating the impact of the channel through which a transmitted signal traverses in a wireless environment. Conventionally, the channel effect is encapsulated in a channel state information (CSI) block in modern communication systems. While conventional methods like Minimum Mean Square Error (MMSE) are employed for CSI estimation, they often entail high computational costs and may not always align with the demands of real networks. Furthermore, obtaining timely CSI information gets more challenging due to extended propagation delays and fast-changing propagation environments in CCS systems [61].
@@
Various techniques have been developed to measure SatCom channels. In [64], Wang et al. introduce a method known as adaptive random-selected multi-beamforming estimation, which focuses on estimating geometric-based millimeter-wave Multiple-Input Multiple-Output (MIMO) CSI. They create a geometry-based channel model for multi-satellite applications within high-throughput satellite systems. A notable feature of this method is its application of compressive sensing, which estimates the combination of the transmit beamformer on the satellite side and the receiving combiner on the user side, organized randomly across several time slots. By exploiting sparsity in the angle domain, this approach significantly reduces the number of measurements needed for precise channel estimation.
The ML-based methods are increasingly being adopted as a promising alternative for channel estimation. This channel estimation can be framed as a supervised learning problem by using various channel features as inputs, including distance, time delay, received power, azimuth angles of arrival (AoA) and departure (AoD), elevation angle, root mean square (RMS) delay spread, and frequency, with CSI serving as the output labels. In [65], the reciprocity property of downlink and uplink channels in time division duplexing (TDD) systems is explored, allowing the downlink channel to be estimated from uplink CSI using an LSTM-based deep learning model. Lu et al. in [66] propose a channel estimation method utilizing a fundamental deep learning architecture in multi-satellite, multi-ground environments.
@@ 3.2 Cloud-Native Distributed MIMO Cooperation and Coordinated Signal Processing
The deployment of large-scale satellite constellations, comprising thousands or even tens of thousands of satellites, has garnered considerable attention from industry and academia. In such systems, multi-satellite MIMO can reduce the reliance on a more significant number of antennas while improving data throughput [67]. One approach utilizes multiple non-collaborating satellites to independently apply MIMO techniques. Alternatively, cooperation between satellites, through the exchange of data or CSI, enables joint tasks such as transmission and resource allocation [68, 69]. This cooperation approach maximizes the benefits of the MIMO technique, significantly enhancing overall system performance. In addition, in CCS system, collaborative reception and transmission among multiple satellite nodes can achieve uplink reception diversity gain and downlink coherent beamforming gain, respectively. In theory, joint transceiver processing can achieve optimal performance. Scalable distributed cooperative baseband signal processing has been extensively studied in cell-free massive MIMO systems [70]. The authors in [71] apply the concept of cell-free massive MIMO to satellite communications, and both theoretical and simulation results demonstrate that this approach can significantly enhance system capacity. However, the cell-free massive MIMO in [71] requires the deployment of a centralized processing unit in space, which is highly challenging to implement.
@@
Based on Option 7 in Figure 4, the authors in [72] propose a new low-layer physical splitting scheme to enable scalable distributed baseband signal processing, further evolving into a cell-free wireless access network. Introducing the new low-layer physical layer splitting scheme into the CCS architecture enables cloud-native collaboration of baseband computing power, thereby reducing the complexity of implementing cooperative transmission. As shown in Figure 8, the Sat-EDU implements distributed baseband transceiver processing, the Sat-DU performs high-layer physical layer processing and upper-layer processing for anchor users, while the Sat-CU handles anchor user selection, user-to-node and beam association, and other control functions.
However, the implementation of distributed MIMO cooperation and on-board coordinated baseband signal processing in CCS system faces significant technical challenges, primarily including the realization of distributed transceivers, synchronization and calibration among nodes. Specifically, when the clock or local oscillator signals are generated locally at each satellite, attaining precise synchronization and calibration in terms of absolute phase, frequency, and time becomes extremely challenging [39]. Also, distributed cooperative baseband signal processing is frequently hampered by the limitations of onboard processing capabilities and inter-satellite interaction capacity. Consequently, there is an urgent need to develop robust, efficient, and resource-conserving synchronization and calibration techniques to bolster the functionality of multi-satellite MIMO systems. In addition to the above challenges, distributed MIMO proves to be of vital importance for the effective management of the vast amounts of data generated by such systems. Through the utilization of decentralized processing nodes, each satellite is enabled to independently process signals and subsequently share the results with neighboring satellites. This approach not only diminishes the overreliance on centralized data processing but also mitigates communication overhead. Moreover, cloud-native architectures take the distributed MIMO cooperation and coordinated baseband signal processing to a new level by furnishing scalable and flexible capabilities that can dynamically adapt to the ever-evolving needs of the network [73].

### v2

@@ 2.3 Reconfigurable Satellite Formation Flying
Moreover, the virtual structure formation method takes a different approach by treating the formation as a single “rigid body” [42]. Each satellite maintains a fixed position within a virtual structure, such as a wheel, line, or grid. This method is effective in achieving high precision in maintaining formation shape and alignment. For example, wheel formations involve satellites distributed around a circular or ring-like path, providing 360-degree coverage ideal for tasks like surveillance and communication relays. Other configurations include linear formations, which arrange satellites in a straight line, suitable for continuous EO and synchronized orbit tracking, and rectangular or grid formations, where satellites form a structured grid pattern for broad-area coverage. Triangular formations and cubic or spherical formations are also common in virtual structure control, offering three-dimensional coverage for complex space missions that require multi-directional observations or high spatial resolution. Additionally, honeycomb formations follow a hexagonal grid structure, maximizing spatial efficiency and are often used in communication networks and environmental monitoring due to their dense, uniform coverage. Arrow or V-formations are another variant, typically used in applications that require directional alignments, such as clustered navigation or coordinated maneuvering, as they reduce aerodynamic drag (for atmospheric applications) and optimize the communication link among satellites. Beyond these conventional configurations, recent research has also introduced optimization and predictive control techniques into satellite formation control. For instance, optimal control methods aim to minimize fuel consumption or time delay, which is essential for deep-space missions with limited resources. Model predictive control (MPC) allows satellites to adjust their trajectories based on predictions of future system states, providing adaptive and efficient control for complex formations with high maneuverability demands.
@@
Several research efforts have focused on developing the necessary technologies to achieve RSFF. One of the most prominent approaches is the use of graph theory to model inter-satellite communication and control [46]. By representing satellites as nodes in a graph and their communication links as edges, researchers have developed algorithms to optimize formation control under various constraints, including limited communication bandwidth, time-varying network topologies, and external disturbances. Another key area of research involves the development of formation control strategies based on artificial potential functions. These methods model satellites as particles moving within a potential field generated by their relative positions and target locations. By adjusting the forces acting on each satellite, the formation can be dynamically reconfigured to achieve the desired geometries while avoiding collisions or other hazards. This approach has proven effective in both simulation and experimental settings, particularly for missions involving close-proximity operations. Moreover, the control methods for swarm configurations share similarities with satellite formation control methods. However, for CCS system already in the desired configuration, frequent control task execution leads to resource wastage. Therefore, an event-triggered mechanism can be designed to apply control only when configuration errors exceed the predefined threshold. Event-triggered control can effectively reduce both communication frequency among members and energy consumption for orbit maintenance. Currently, event-triggered control has been increasingly utilized for maintaining swarm system configurations, specifically in research areas such as first-/second-order integral multi-agent systems [47], directed/undirected topologies [48], centralized/distributed triggering mechanisms [49], fixed/switched topologies [47], linear/nonlinear systems [50], with/without time delays [51], and input saturation [52].
@@
Further, developing decentralized control architectures that distribute decision-making across the DSIN is essential to improve resilience and flexibility. The DSIN should incorporate adaptive algorithms capable of responding to environmental changes and communication delays, enabling autonomous reconfiguration without relying on a central command unit. Moreover, hybrid centralized and decentralized control architectures use centralized control for high-precision tasks and decentralized strategies for adaptive and resilient reconfiguration, enabling the balance between precision and robustness. For inter-satellite distances exceeding 20 km, orbit maneuver requirements can be addressed based on the Lambert problem. The Lambert problem is a typical two-point boundary value problem that determines the maneuver trajectory given the spacecraft’s initial and final positions and the transfer time [53]. Considering the univariate nature of the Lambert problem, recent studies have derived transfer time equations using parameters such as flight path angle [54], eccentricity [55], and terminal velocity [56], and proposed iterative solution methods. Furthermore, maneuver theory has been improved by incorporating factors such as environmental perturbations [57], thrust direction constraints [58], and multi-revolution orbit solutions [59]. The last but not least, researching algorithms that optimize energy consumption during formation reconfiguration is also crucial. These algorithms could focus on minimizing fuel use while considering the long-term sustainability of the mission. Techniques such as cooperative energy management and predictive energy budgeting can be explored to extend the operational lifespan of satellite constellations.
@@
By adopting a phased configuration control method at different spatial scales, the CCS systems can be restructured according to service demands, choosing system configurations and control maintenance strategies that best fit the task requirements. This enables the design of highly flexible, scalable, and open network architectures for DSIN, along with the corresponding functional partitioning strategies. The future of RSFF lies in the continued integration of advanced control algorithms, high-precision sensors, and scalable communication networks. One promising avenue is the application of machine learning techniques to predict and adapt to environmental disturbances in real time. By incorporating predictive models into the control loop, satellite formations can proactively adjust their configurations to maintain optimal performance. Another important direction is adopting a phased configuration control method at different spatial scales, which allows the CCS system to be restructured according to service demands and to choose system configurations and control maintenance strategies that best fit the task requirements.
@@ 3 Enabling technologies
In order to achieve spectrum- and energy-efficient communication, as well as high timeliness and reliability, a variety of new air interface and transmission technologies will be used in the DSIN, especially in the CCS system. In this section, we will first introduce the air interface technologies at the physical and link layers, including channel modeling and estimation, new waveforms, distributed antennas, channel coding methods, multiple access approaches, and multicasting mechanisms, followed by network and transport layer technologies, involving erasure transfer protocols, distributed routing, and congestion control.
@@ 3.1 Channel Modeling and Estimation for CCS System
As the dawn of 6G communication technologies beckons, the crucial role of DSIN in achieving uninterrupted global connectivity has become more pronounced than ever [60]. Within this evolving landscape, the creation of precise channel models and the enhancement of channel estimation methods have become pivotal to the advancement of these systems. Channel models are indispensable for capturing the intricacies of satellite-ground communication links, offering a fundamental insight into their performance across diverse operational scenarios. Concurrently, channel estimation stands as a cornerstone technology, facilitating the real-time acquisition of channel state information, which is indispensable for fine-tuning signal transmission and guaranteeing the efficiency and dependability of DSIN [61]. Despite extensive research into terrestrial communication channel modeling and estimation, the unique challenges posed by satellite-to-ground channels in CCS system demand a closer look. In this subsection, we delve into the channel modeling and estimation within the realm of CCS system, and explore the latest advancements and present a comprehensive review of the current research landscape.
@@ 3.1.1 Channel Model
In CCS system, the channel model is essential to understand and predict the behavior of signals propagating between satellite-to-satellite and satellite-to-user links. We delve into the two primary categories of the channel model, the satellite-to-satellite and user-satellite channel model, focusing on the user-satellite channel model [62].
The satellite-to-satellite channel model, which concerns the communication links between satellites, is relatively straightforward compared to user-satellite modeling. This simplicity arises because space-to-space communication does not involve the complexities of the Earth’s atmosphere or terrestrial obstacles. As shown in Figure 7, in satellite-to-satellite channel models, the primary considerations include the relative positions and movements of the satellites, the distances between them, and the characteristics of the transmission medium, which is typically the vacuum of space. The main challenges in the satellite-to-satellite channel model involve accounting for the dynamic nature of CCS system, including the effects of gravitational forces, satellite orbit perturbations, and the need for precise alignment of communication beams. However, the relative ease of this model allows researchers to focus on optimizing the performance of ISL without the added complexities of atmospheric and terrestrial interference. The user-satellite channel model is where the real challenge lies. This type of model ensures reliable communication between satellites and ground stations [63]. The user-satellite channel model is more challenging than the satellite-satellite channel model, primarily due to many factors, including the influence of the Earth’s atmosphere, interference from terrestrial obstacles, and the dynamic nature of the communication environment. User-satellite channel model methods mainly include geometric stochastic models and machine learning (ML)-based models.
@@
The geometric stochastic model is designed with rational assumptions regarding the distribution of obstacles and scatterers near the receiving user. It abstracts the effective scatterers within the environment into one or more models characterized by distinct geometric forms, systematically distributing them across the environment’s abstracted geometric structure. In [64], the studies undertook the channel model of a multi-satellite communication system, postulating that the satellites are uniformly dispersed across the surface of a sphere at a specified altitude above the Earth. Using the geometric relationship between the satellite and the Earth, the research deduced the correlations among the elevation angle, azimuthal angle, and the distance between the satellite and the Earth. Although the geometric stochastic model has high universality and low computational complexity, it has low modeling accuracy in specific scenarios. To have a more accurate channel model of CCS system, much research is needed to consider more factors.
In CCS wireless communication environments, the channel conditions are more complex, the operating frequency bands are higher, and the mobility of the terminals is further enhanced. In such scenarios, ML’s self-learning and predictive capabilities are particularly suitable, as they can extract critical features from highly complex channel data. Moreover, by thoroughly training the channel model, we can enable the model to better adapt to the rapid changes in channel states in high-dynamic scenarios, thereby significantly improving the generalization ability and performance of the model. The authors in [65] integrate ML with traditional channel modeling techniques, presenting a novel approach for channel modeling in LEO satellite systems. It underscores the critical role of radio channel forecasting in improving link quality in the face of growing atmospheric impairments. In recent years, machine learning-based satellite-Earth channel modeling has seen rapid advancement, with many models being proposed, showcasing the technology’s robust learning capacity and proficiency in handling vast datasets. The long short-term memory (LSTM)-based modeling has garnered significant attention, with superior prediction accuracy demonstrated in several studies. However, the predictive accuracy of current ML-based models is contingent upon the richness of the training dataset and demands substantial computational resources, which limits the applicability of many models to specific scenarios or frequency bands, as noted in [66].
@@ 3.1.2 Channel Estimation
Channel estimation is an essential aspect of CCS systems, and is a critical issue that directly affects the performance and reliability of the communication link. This entails estimating the impact of the channel through which a transmitted signal traverses in a wireless environment. Conventionally, the channel effect is encapsulated in a channel state information (CSI) block in modern communication systems. While conventional methods like Minimum Mean Square Error (MMSE) are employed for CSI estimation, they often entail high computational costs and may not always align with the demands of real networks. Furthermore, obtaining timely CSI information gets more challenging due to extended propagation delays and fast-changing propagation environments in CCS systems [67].
@@
Various techniques have been developed to measure SatCom channels. In [70], Wang et al. introduce a method known as adaptive random-selected multi-beamforming estimation, which focuses on estimating geometric-based millimeter-wave Multiple-Input Multiple-Output (MIMO) CSI. They create a geometry-based channel model for multi-satellite applications within high-throughput satellite systems. A notable feature of this method is its application of compressive sensing, which estimates the combination of the transmit beamformer on the satellite side and the receiving combiner on the user side, organized randomly across several time slots. Taking advantage of the sparsity in the angle domain, this approach significantly reduces the number of measurements needed for precise channel estimation.
ML-based methods are increasingly being adopted as a promising alternative for channel estimation. This channel estimation can be framed as a supervised learning problem by using various channel features as inputs, including distance, time delay, received power, azimuth angles of arrival (AoA) and departure (AoD), elevation angle, root mean square (RMS) delay spread, and frequency, with CSI serving as the output labels. In [71], the reciprocity property of downlink and uplink channels in time division duplexing (TDD) systems is explored, allowing the downlink channel to be estimated from uplink CSI using an LSTM-based deep learning model. Lu et al. in [72] propose a channel estimation method utilizing a fundamental deep learning architecture in multi-satellite, multi-ground environments. In [73], a ML-based CSI prediction method is applied in a massive MIMO communication scenario, using convolutional neural networks to extract temporal channel correlation features. In [74], an LSTM-based predictor is used to address the problem of channel ageing in LEO satellite communication systems. Meanwhile, several DL-based joint CSI prediction methods are incorporated in downlink precoding schemes[75, 76]. A DL-based quantized phase hybrid precoder is introduced to enhance spectral efficiency in [75]. Furthermore, taking channel prediction errors into account, [76] proposes a deep learning (DL)-based joint channel prediction and multibeam precoding scheme, which achieves significant gains in robust CSI acquisition and uplink transmission performance, even under high Doppler shifts and long propagation delays.
Although AI-enabled channel estimation methods provide a favourable trade-off between generalisation and performance-complexity, they lack interpretability in decision-making. To address this issue, [77] proposes a novel explainable AI (XAI)-based channel estimation scheme to offer detailed and reasonable interpretability of deep learning (DL) models. In summary, AI-enabled channel estimation methods have demonstrated unique advantages in enhancing the reliability and robustness of satellite-terrestrial communications. However, the significant mismatch between the learning capabilities of existing AI models and the constraints of onboard satellite systems remains a challenging issue. Therefore, guiding AI-enabled channel estimation methods toward lightweight and universally reconfigurable solutions will be a key focus for future CCS system.
@@ 3.2 Cloud-Native Distributed MIMO Cooperation and Coordinated Signal Processing
The deployment of large-scale satellite constellations, comprising thousands or even tens of thousands of satellites, has garnered considerable attention from industry and academia. In such systems, multi-satellite MIMO can reduce the reliance on a more significant number of antennas while improving data throughput [78]. One approach utilizes multiple non-collaborating satellites to independently apply MIMO techniques. Alternatively, cooperation between satellites, through data exchange or CSI, enables joint tasks such as transmission and resource allocation [79, 80]. This cooperation approach maximizes the benefits of the MIMO technique, significantly enhancing overall system performance. In addition, in CCS system, collaborative reception and transmission among multiple satellite nodes can achieve uplink reception diversity gain and downlink coherent beamforming gain, respectively. In theory, joint transceiver processing can achieve optimal performance. Scalable distributed cooperative baseband signal processing has been extensively studied in cell-free massive MIMO systems [81]. The authors in [82] apply the concept of cell-free massive MIMO to satellite communications, and both theoretical and simulation results demonstrate that this approach can significantly enhance system capacity. However, the cell-free massive MIMO in [82] requires the deployment of a centralized processing unit in space, which is highly challenging to implement.
@@
Based on Option 7 in Figure 4, the authors in [83] propose a new low-layer physical splitting scheme to enable scalable distributed baseband signal processing, further evolving into a cell-free wireless access network. Introducing the new low-layer physical layer splitting scheme into the CCS architecture enables cloud-native collaboration of baseband computing power, thereby reducing the complexity of implementing cooperative transmission. As shown in Figure 8, the edge distributed units (Sat-EDU) implements distributed baseband transceiver processing, the Sat-DU performs high-layer physical layer processing and upper-layer processing for anchor users, while the Sat-CU handles anchor user selection, user-to-node and beam association, and other control functions.
However, the implementation of distributed MIMO cooperation and on-board coordinated baseband signal processing in CCS system faces significant technical challenges, primarily involving the realization of distributed transceivers, as well as synchronization and calibration among nodes. Specifically, when the clock or local oscillator signals are generated locally at each satellite, attaining precise synchronization and calibration in terms of absolute phase, frequency, and time becomes extremely challenging [45]. Also, distributed cooperative baseband signal processing is frequently hampered by the limitations of on-board processing capabilities and inter-satellite interaction capacity. Consequently, there is an urgent need to develop robust, efficient, and resource-conserving synchronization and calibration techniques to bolster the functionality of multi-satellite MIMO systems. In addition to the above challenges, distributed MIMO proves to be of vital importance for the effective management of the vast amounts of data generated by such systems. Through the utilization of decentralized processing nodes, each satellite is enabled to independently process signals and subsequently share the results with neighboring satellites. This approach not only diminishes the overreliance on centralized data processing, but also mitigates communication overhead. Moreover, cloud-native architectures take the distributed MIMO cooperation and coordinated baseband signal processing to a new level by furnishing scalable and flexible capabilities that can dynamically adapt to the ever-evolving needs of the network [84].